Quick Question: What has happened to the purchasing power of the minimum wage over this period (i.e., has it increased or decreased)? And the % change between 1965 and 2019 for each of the columns. Thanks!

Year min wage CPI 1965 = 100 Real wages based on 1965 prices
1965 1 100 1
1967 2.65 185.12 1.43
2019 14 809.52 1.73
Answer #1

The purchasing power of the minimum wage has increased because real wages have increased from 165 to 2019.
% change between 1965 and 1967 = [(1.43-1)/1]*100 = (0.43)*100 = 43%
% change between 1967 and 2019 = [(1.73-1.43)/1.43]*100 = (0.3/1.43)*100 = 20.98%